OpenStudy (anonymous):

Two similar hexagons have areas 36 sq. in. and 64 sq. in. The ratio of a pair of corresponding sides is 9/16. True or False?

OpenStudy (mathmale):

Hello, and welcome to OpenStudy! 0 Have you considered sketching ing the two hexagons, or figuring out the areas of each? Note that "hexa=" implies 6 "pie slices." Every central angle is 360 deg / 6 = 60 deg. This info should enable you to find a formula for the area of each hexagon. Compare the radii of the two hexagons you've drawn, by writing a ratio. Is your ratio equal to or equivalent to 9/16?

OpenStudy (anonymous):

 then take the ratio of areas

jhonyy9 (jhonyy9):

36 9 --- = --- 64 16 4*9 9 ----- = --- 4*16 16 9 9 --- = --- hope this make sens giving the right answer easy 16 16

Directrix (directrix):

If two polygons are similar, the square of the scale factor of the two polygons is equal to the ratio of the areas of the polygons.

Directrix (directrix):

Is (9 / 16) ^2 = ( 36 / 64 ) ?